OpenStudy (anonymous):

4/3 + 1/5 = 1/x solve

OpenStudy (anonymous):

23/15 = 1/x x = 15/23

OpenStudy (anonymous):

\[4/3 +1/5=1/x \] \[20/15 + 3/15=1/x\] \[23/15=1/x\], so flipping the expressions on both sides gives you x=15/23

OpenStudy (anonymous):

thank you. I am still a little confused on how you got those numbers